OpenStudy (anonymous):

solve by the elimination method 5x+6y=5 10x+12y=10

OpenStudy (anonymous):

that's the same equation twice - you sure there's no typo?

OpenStudy (anonymous):

no solution

OpenStudy (anonymous):

That is the exact way the problem is written

OpenStudy (anonymous):

-2(5x+6y=5) = -10x-12y=-10 10x+12y= 10 they all eliminate meaning no solution

OpenStudy (anonymous):

THANKS GUYS

OpenStudy (anonymous):

actually - that's wrong. For example, one particular solution is x = 1, y = 0 ...

OpenStudy (anonymous):

that is true but it asks to solve by the elimination method, can you get that answer solving it using that particular method.

OpenStudy (anonymous):

hmm - you can't use the elimination method on a single equation in the first place, or can you? So the assignment is kinda.. ahm.. yeah - you got to figure put what your teacher wants to hear from you, i guess :P

OpenStudy (anonymous):

then again.. with one equation in 2 variables, it's no intersection-POINT ;)

OpenStudy (anonymous):

If that's true then the solution is all points on the line

OpenStudy (anonymous):

actually, they both intersect at all points on a line

OpenStudy (anonymous):

that's right... the "slope-intercept-form" is:\[y = -{5 \over 6} x + {5 \over 6}\]